A cubical block of wood of length 10cm floats at the interface between oil of density 800kg/m^(3) and water.The lower surface of the block is 1.5 cm below the interface.If the depth of water is 10cm below the interface and oil is up to 10 cm above the interface then the difference in pressure at the lower and the uppper face of the wooden block is (Assume density of water =1000kg/m^(3) and g=10m/s^(2)
1
Expert's answer
2020-01-22T04:24:07-0500

The gauge pressure at the top is


"p_2-p_0=\rho_ogh=(800)(9.8)(0.015)=118\ Pa"

The gauge pressure at the bottom is

"p_1-p_0=\rho_ogh_o+\rho_wgh_w"

"p_1-p_0=(800)(9.8)(0.1)+(1000)(9.8)(0.015)=931\ Pa"

The difference in pressure at the lower and the uppper face of the wooden block is


"p_1-p_2=931-118=813\ Pa"